OpenStudy (anonymous):

Find the constant sum for an ellipse with foci F1 (-5, 0), F2 (5, 0) and the point on the ellipse (0, -12). The constant sum is: _____. Enter only a number.

OpenStudy (anonymous):

23

OpenStudy (anonymous):

here are the big guns ;)

OpenStudy (amistre64):

the sum of the lengths of the lines from foci to point is the constant sum; since an ellipse is defined as the locus of points that are a constant sum from the foci

OpenStudy (amistre64):

they was nice enough to even put everything on one axis for ya :)

OpenStudy (amistre64):

well, then again ; i think a transposed some coords

OpenStudy (anonymous):

constant sum= sum of the distances between any point on the ellipse to the two foci constant sum=sqrt((0+5)^2+(-12-0)^2)+sqrt((0-5)^2+(-12-0)^2) =sqrt(25+144)+sqrt(25+144) =sqrt(169)+sqrt(169) =13+13 =26

OpenStudy (amistre64):



OpenStudy (anonymous):

wow, really thorough answers guys! thankyou i actually get it now